radomes.openskiesproject.org
k5bookRenCasillme
k5bookRenCasillme k5bookRenCasillme
  • 17-08-2016
  • Mathematics
Answered

a game has 40 chips, each labeled with a number from 1 through 40. ben randomly chooses one chip. what is the probability that he chooses 10 or a multiple of 8? a.1/8 b.3/20 c.1/6 d.1/320

Verified Answer :

texaschic101
texaschic101 texaschic101
  • 17-08-2016

chooses 10 or a multiple of 8/ total chips

choosing a 10.......1/40
choosing a multiple of 8.....5/40

1/40 + 5/40 = 6/40 = 3/20 <===

Answer Link

Related Questions

Pls help me guys you guys are talented
How do you diagram the sentence The beautiful white swans floated
Which of the following is a conservative viewpoint A Large government encourages economic growth and individual choice B Large government endangers economic gro
What are the effects of human activity
Toms home is 1800 m from his school One morning he walked part of the way to school and then ran the rest If it took him 20 minutes or less to get to school and
Mark had been a resident of Larchmont for 17 years and local folks thought of him as a solid citizen In this sentence the word solid probably means A heavy B d
PLEASE HELP ME What would be an immediate effect of the construction of roads and parking lots on the transfer of energy from producers to consumers in an ecosy
Can someone please help me on the second quick Check
Order 34x102 12x107 811x103 and 435 from least to greatest
Describe the lifestyle of English planters in the New World